OpenStudy (anonymous):

What multiplies to get -81 but adds to get -24

OpenStudy (aaliyahfairgood):

-27 times 3 = -81 -27+3=-24

OpenStudy (anonymous):

x*y=-81 x+y=-24 find x and y

OpenStudy (anonymous):

-27 and 3

OpenStudy (anonymous):

this is guess work only,but one i mentioned is general \[y=\frac{ -81 }{ x }\] \[x-\frac{ 81 }{ x }=-24\] \[x^2-81=-24x\] \[x^2+24x-81=0\] factorize and find the values of x and then corresponding value of y